40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Entry  Tue Mar 22 16:29:04 2016, tara, DailyProgress, PMC, Epoxy glue on PMC mirror and PZT 
    Reply  Thu Mar 24 19:07:38 2016, tara, DailyProgress, PMC, Epoxy glue on PMC mirror and PZT 
       Reply  Sat Mar 26 13:36:09 2016, tara, DailyProgress, PMC, Epoxy glue on PMC mirror and PZT IMG_2468.JPGIMG_2466.JPG
          Reply  Tue Mar 29 23:23:37 2016, rana, DailyProgress, PMC, Epoxy glue on PMC mirror and PZT 
             Reply  Wed Mar 30 08:29:15 2016, tara, DailyProgress, PMC, test on new PMC IMG_2528.JPGIMG_2540.JPG
                Reply  Wed Mar 30 10:28:44 2016, Evan, DailyProgress, PMC, test on new PMC 
                   Reply  Thu Mar 31 15:11:27 2016, tara, DailyProgress, PMC, test on new PMC 
                      Reply  Tue Apr 5 16:32:36 2016, tara, DailyProgress, PMC, test on new PMC 
                         Reply  Wed Apr 6 15:02:50 2016, tara, DailyProgress, PMC, test on new PMC 
                            Reply  Fri Apr 8 09:44:49 2016, tara, DailyProgress, PMC, test on new PMC IMG_2634.JPGIMG_2635.JPGIMG_2633.JPG
Message ID: 1620     Entry time: Thu Mar 31 15:11:27 2016     In reply to: 1619     Reply to this: 1623
Author: tara 
Type: DailyProgress 
Category: PMC 
Subject: test on new PMC 

The measured finesse for the new PMC is ~ 300. The previous calculation was wrong because of the wrong calibration.

 

Finesse measurement revisited I rechecked the Finesse measurement and found that I got the same result, so I realized that the calibration might be wrong. So here is the explanation

  • The calibration of 3 MHz/V is wrong.  The measurement is reported in PSL:182. It is for the laser in the South path. The laser has been in the original setup since 2009. The more recent calibration (4.4 MHz/V) wasd done by Evan,  see PSL:1478.
  • The calibration of 4.5 MHz/V is for laser in the North path, which is the path we use in the PMC test. I wrote down the calibration in PSL:1393.
  • The correct calibration with the measurement yields the Finesse of 308 +/-7 (the design is 310).  It seemed our assembly process was quite ok, no serious extra loss observed. The error is mostly from the uncertainty in the FWHM measurement ( 0.90 +/- 0.01 mV), and the error in the calibration (say 4.5 +/- 0.1 [MHz/V]). The error in FSR is smaller ~ 1249 +/- 1 MHz, assuming that the epoxy changes the mirror position by 0.1 mm, so it does not contribute much. 

PZT range revisited:   I mentioned in the previous entry that I can find several (4-5) TEM00 within 0-100 V applied to the PZT. This number should be wrong because we can expect only ~ 6 resonance over 200 V (3 um), or ~ 3 resonance over 100 V. It turned out that what I saw was not real TEM 00, but something else that made me think I saw some hysteresis.

  • The PZT range is 3.3 um with 200 V input, so it is roughly 16 nm/V. We should be able to scan over 1 FSR if the PZT stretch out by lambda/2. And this is what I found. TEM00 occured at every 30V from 0 - 100 V, which corresponded to 30 V* 16 nm/V =0.5 um (lambda/2). So PZT is quite linear and works well.
  • The problem is, between each resonance I saw other resonance modes  1) Another TEM00, but with ~ 6 V apart from the main TEM00. The shape was similar to TEM 00, but the transmitted power is ~ 20 % less. 2) I also saw are other TEM00 like modes that were quite bright, but their powers were negligible (less than a few percent of the main TEM00). It appearred and went away fast, I could not measure the corresponding voltage carefully. These unexpected resonance made me think it was hysteresis in PZT. 
  • For the TEM00 with 5 V apart (~ 80 nm or ~ 10 MHz), I thought it was from some polarization effect. I changed the polarization using half wave plate, but I did'nt see any change in the power ratio of the two modes.
  • Antonio suggested that it might be from the internal back reflection of the mirror because both surfaces are parallel. I'll need to think of a way to check that.

To Do Next: Anyway, the PMC/PZT are good. We will try to lock it tomorrow.

  • Need to find the EOM base to install an EOM from PMC in the North path ( I remembered I had a spare but couldn't find it).
  • Turn on the machine for controlling PMC card. (Need to ask Aidan how to do that).

 

 

ELOG V3.1.3-